OpenStudy (anonymous):

Find the variation constant and an equation of variation where y varies inversely as x and y=11 when x=6.

OpenStudy (callisto):

\[y \alpha 1/x\] y = k(1/x) sub y=11 when x=6 into the equation 11 = k(1/6) k =11x6 = 66